OpenStudy (snowystorm777):

Simplify the following expression: 2x − 4y + 3x2 + 7y − 16x

OpenStudy (snowystorm777):

help fan and medal!

OpenStudy (gabbyalicorn):

2x−4y+3x2+7y−16x =2x+−4y+3x2+7y+−16x Combine Like Terms =2x+−4y+3x^2+7y+−16x =(3x^2)+(2x+−16x)+(−4y+7y) =3x^2+−14x+3y Answer: =3x2−14x+3y